OpenStudy (anonymous):

whats the product of sqrt 3 (2b sqrt49 -b)

OpenStudy (mandre):

\[2 \sqrt{3}\sqrt{49}-\sqrt{3}b\]

OpenStudy (anonymous):

how do i do that. Im clueless on this subject

OpenStudy (mandre):

I missed a b in the first one \[2 \sqrt{3}(-7)b-\sqrt{3}b\] or \[2 \sqrt{3}(7)b-\sqrt{3}b\] Just multiply each term with sqrt 3

OpenStudy (anonymous):

ohhh so would the answer be 13b sqrt 3?

OpenStudy (mandre):

It has 2 answers because sqrt 49 can be -7 or 7.

OpenStudy (anonymous):

My options are bthe square root of 3. 13bthe square root of 3. 7b 7bthe square root of 3.

OpenStudy (mandre):

That would be 2 * 7 - 1 (sqrt 3) The -1 is for the last term.

OpenStudy (mandre):

They take it as only +7. (2 * 7 - 1)* (b) * (sqrt 3) would be your answer.

OpenStudy (anonymous):

Thats not one of my options

OpenStudy (mandre):

LOL What is (2 * 7 - 1). Simple math :)

OpenStudy (anonymous):

13?

OpenStudy (mandre):

Yup

OpenStudy (anonymous):

oh i didnt see that option, sorry. But thank you so much!

OpenStudy (mandre):

Pleasure.